OpenStudy (anonymous):

mulitiply simpliest form 9x/56mulitiply7/3x

OpenStudy (pfenn1):

Do you mean \[\left( 9x \over 56 \right) \times \left( 7 \over 3x \right)\]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

You can simplify the expressions because 3 goes into 9, the x cancels, and 56 is divisible by 7.

Parth (parthkohli):

= 3x/8

Parth (parthkohli):

By cancelling the cross expressions, you get 3x/8 * 1/1 which is obviously 3x / 8

OpenStudy (pfenn1):

@ParthKohli, the x's cancel out as well.

Parth (parthkohli):

Alright, thanks, so it is 3/8 (it can also be 3x/8 but not the simplest form)

OpenStudy (pfenn1):

no, it could also be 3x/8x but not 3x/8